OpenStudy (anonymous):

The steeple on a building is a square pyramid with a base area of 12 square feet and a height of 15 feet. How many cubic feet of concrete was used to make the steeple? Help!

OpenStudy (anonymous):

you do 12 times 15 and you get ...... 180.

OpenStudy (hoblos):

volume of pyramid = Ah/3 where A=area of base , h= height so here V= 12*15/3 = 180/3 = 60

OpenStudy (anonymous):

thank you.